# The global attractor of the inelastic linear Boltzmann equation in a gravity field for Maxwell molecules

### v1

@@ 3.1 Moments estimates
From the weak formulation (1.28) it is possible to derive a system of evolution equations for $M_{0},M_{1},M_{2}$ by choosing as test function $\left\{1,v,|v|^{2}\right\}$ respectively. It is important to note that the functions $\{1,v,|v|^{2}\}$ are not test functions in the sense of Definition 1 and we should consider suitable cut-offs of the test functions. However, the result obtained in the by following this procedure is the same as the one we would obtain formally choosing $\{1,v,|v|^{2}\}$ as test functions. We refer, for instance, to Remark 6 in [56] for a detailed discussion on this technical issue. Therefore, to keep the exposition as simple as possible we will take $\{1,v,|v|^{2}\}$ as test functions in (1.28). Hence, using also the collision rule (1.8) leads to
$\begin{cases}\vspace{2mm}\displaystyle{\frac{\mathop{}\!\mathrm{d}M_{0}(t)}{\mathop{}\!\mathrm{d}t}=0},&\\ \vspace{2mm}\displaystyle{\frac{\mathop{}\!\mathrm{d}M_{1}(t)}{\mathop{}\!\mathrm{d}t}=aM_{0}(t)-(1+r)\int_{\mathbb{R}^{d}}f(t,dv)\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(v\cdot\omega)\omega\mathop{}\!\mathrm{d}\omega},&\\ \displaystyle{\frac{\mathop{}\!\mathrm{d}M_{2}(t)}{\mathop{}\!\mathrm{d}t}=2a\cdot M_{1}(t)-2(1-r^{2})\int_{\mathbb{R}^{d}}f(t,dv)\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(v\cdot\omega)^{2}\mathop{}\!\mathrm{d}\omega}.\end{cases}$ (3.1)
The previous system can be integrated exactly. This is the purpose of the next propositions.
@@
Let $d=2$, let $a\in\mathbb{R}^{d}$ and let the restitution coefficient $0<r\leq 1$. Let $f\in C([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ the unique weak solution to (1.26) given by Theorem 1.1. Suppose $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$, then we have that $M_{0}(t),M_{1}(t),M_{2}(t)\in L^{\infty}([0,+\infty))$.
@@
We start from (3.1) and we consider a rotation $R$ such that $n=Re_{1}$ and $\omega=R\tilde{\omega}$. For $d=2$ we have $\tilde{\omega}=(\cos\theta,\sin\theta)$ with $\theta\in[0,2\pi]$. Therefore we have:
@@
Therefore for we obtain the following
@@
$M_{0}(t)$ $=M_{0}(0);$ (3.10)
$M_{1}(t)$ $=e^{-C_{1}t}M_{1}(0)+\frac{aM_{0}(0)}{C_{1}}\left(1-e^{-C_{1}t}\right);$ (3.11)
@@
Let $d=3$, let $a\in\mathbb{R}^{d}$ and let the restitution coefficient $0<r\leq 1$. Let $f\in C([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ the unique weak solution to (1.26) given by Theorem 1.1. Suppose $\{|v|,|v^{2}|\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$, then we have that $M_{0}(t),M_{1}(t),M_{2}(t)\in L^{\infty}([0,+\infty))$.
@@
$-(1+r)\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(v\cdot\omega)\omega\mathop{}\!\mathrm{d}\omega$ $=-2\pi(1+r)\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)v;$ (3.13)
$-(1-r^{2})\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(n\cdot\omega)^{2}\mathop{}\!\mathrm{d}\omega$ $=-2\pi(1-r^{2})\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)|v|^{2}.$ (3.14)
@@
$M_{0}(t)$ $=M_{0}(0);$ (3.15)
$\frac{\mathop{}\!\mathrm{d}M_{1}(t)}{\mathop{}\!\mathrm{d}t}$ $=aM_{0}(0)-2\pi(1+r)\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)M_{1}(t);$ (3.16)
@@
Set $\tilde{C}_{1}=2\pi(1+r)\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right),\tilde{C}_{2}=2\pi(1-r^{2})\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)$ for brevity and notice that $\tilde{C}_{1}-\tilde{C}_{2}>0$. Integrating we get
$M_{0}(t)$ $=M_{0}(0);$ (3.18)
$M_{1}(t)$ $=e^{-\tilde{C}_{1}t}M_{1}(0)+\frac{aM_{0}(0)}{\tilde{C}_{1}}\left(1-e^{-\tilde{C}_{1}t}\right);$ (3.19)
@@ 3.2 Existence of a steady state
To prove existence of a steady state we will use the result of Proposition 3.1 and Proposition 3.2 which allows us to use Schauder fixed point theorem to prove the existence of a stationary solution to (1.11) in the sense of Definition 2. In fact, without loss of generality choose $M_{0}(0)=1$ in (3.10)-(3.18) and define
$\mathscr{U}=\left\{f\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)\,:\,\int_{\mathbb{R}^{d}}f(\mathop{}\!\mathrm{d}v)=1,\int_{\mathbb{R}^{d}}|v|f(\mathop{}\!\mathrm{d}v)\leq aC_{0},\int_{\mathbb{R}^{d}}|v|^{2}f(\mathop{}\!\mathrm{d}v)\leq|a|^{2}\hat{C}_{0}\right\}$ (3.21)
@@
Let $d=2,3$, let $a\in\mathbb{R}^{d}$ and let the restitution coefficient be $10<r<$ 1. Let $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ such that $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ and let $f\in\mathcal{C}([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ be the unique weak solution obtained in Theorem 1.1. Then there exists a stationary solution $f_{\infty}$ of (1.28) in the sense of Definition 2 or equivalently a fixed point for the adjoint $S^{*}(t)$ of the semigroup $S(t)$ defined in (3.24).
@@
Due to Proposition 3.1 and Proposition 3.2 the set $\mathscr{U}$ is convex and closed in the weak* topology of $\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. In addition, $\mathcal{U}$ is metrizable and hence sequentially compact and $\mathscr{U}$ is also weak* compact. To see this notice that $\mathscr{U}$ is contained in the unit ball of $\mathscr{M}(\mathbb{R}^{d})$ and the space $\mathcal{C}_{0}(\mathbb{R}^{d})$ is separable $d=2,3$. Since $\mathscr{U}$ is weak*-closed by the Banach-Alaoglu’s theorem (cf. [21]) it follows that $\mathscr{U}$ is compact. We have that for any $h\geq 0$ $S^{*}(h)\mathscr{U}\subseteq\mathscr{U}$, hence the operator $S^{*}(h)$ is weak* compact. Therefore we can apply Schauder fixed point theorem to prove the existence of $f_{*}^{(h)}$ such that $S^{*}(h)f_{*}^{(h)}=f_{*}^{(h)}$ and by the semigroup property $S^{*}(mh)f_{*}^{(h)}=f_{*}^{(h)}$ for every $m\in\mathbb{N}$. Take a sequence $\{h_{k}\}_{k}$ such that $h_{k}\rightarrow 0$ and its corresponding sequence of fixed points $\left\{f_{*}^{(h_{k})}\right\}_{k}$. This sequence is compact in $\mathscr{U}$, since $\mathscr{U}$ itself is compact, and taking a sub-sequence if needed we have $f_{*}^{(h_{k})}\rightarrow f_{*}$ for some $f_{*}\in\mathscr{U}$ as $k\rightarrow+\infty$. For any $t>0$ there exists a sequence of integers such that $m_{k}h_{k}\rightarrow t$ as $k\rightarrow+\infty$. On one hand this yields $S^{*}(m_{k}h_{k})f_{*}^{(h_{k})}=S^{*}(h_{k})f_{*}^{(h_{k})}=f_{*}^{(h_{k})}\rightarrow f_{*}$ while on the other hand
@@
By the weak* continuity of the semigroup $S^{*}(t)$ it follow that the right-hand side converges to $S^{*}(t)f_{*}$ which gives
@@
for every $t\geq 0$. Therefore $f_{*}$ is a fixed point for $S^{*}(t)$ and the proof is concluded. ∎
3.3 Uniqueness
In this subsection we prove the uniqueness of the stationary solution obtained in Theorem 3.3. We will prove uniqueness expressing the stationary solution $f_{\infty}$ by expanding the equation $f_{\infty}=S^{*}(t)f_{\infty}$ by means of a Neumann series which we then prove to be convergent. This is a suitable adaptation of a well-known strategy, see e.g. [4]. The main theorem of this subsection is the following
@@ 3.3 Uniqueness
Let $d=2,3$, let $a\in\mathbb{R}^{d}$ and let the restitution coefficient be $10<r<$ 1. Let $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ such that $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Then the stationary solution obtained in Theorem 3.3 is unique for every $t>0$.
@@
The semigroup $S(t)$ introduced in Theorem 1.1 acts on $\mathcal{C}_{0}(\mathbb{R}^{d})$ as
@@
where $F(t)$ is the semigroup generated by the operator $S\varphi=-a\cdot\partial_{v}\varphi-\nu\varphi$ with domain $\mathscr{D}(T)$. The explicit action of $F(t)$ on $\mathcal{C}_{0}(\mathbb{R}^{d})$ is given by
$F(t)f=e^{-\nu t}f(v-at).$ (3.25)
@@
To compute the adjoint of $V(S(t))$ take $f\in\mathscr{M}(\mathbb{R}^{d}),\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$, we have that
@@
[…]
@@
where the series is absolutely convergent for any $t>0$ and therefore $R(1,F^{*}(t))$. Hence (3.30) reads
@@
$\leq e^{-\nu nt}\sup_{t\geq 0}\|S(t)\|\|\mathscr{L}_{+}\|\int_{0}^{t}\|F^{*}(t-s)\|\mathop{}\!\mathrm{d}s$
$\leq e^{-\nu nt}\nu\int_{0}^{t}e^{-\nu(t-s)}\mathop{}\!\mathrm{d}s=e^{-\nu nt}(1-e^{-\nu t}).$ (3.33)
where we used that fact that $\|S(t)\|\leq 1$ for every $t\geq 0$ since $S(t)$ is a Markov semigroup and that $\|\mathscr{L}_{+}\|\leq\nu$. This give
$\Bigg{\|}\sum_{n=0}^{+\infty}(F^{*}(t))^{n}(V(S(t)))^{*}\Bigg{\|}$ $\leq(1-e^{-\nu t})\sum_{n=0}^{+\infty}e^{-n\nu t}=(1-e^{-\nu t})\frac{1}{1-e^{-\nu t}}$ (3.34)
for any $t>0$ and this shows that the series is convergent. Since the Neumann series (3.32) identifies a unique element we conclude that the stationary solution $f_{\infty}$ is unique, see for instance [4]. ∎

### v2

@@ 3.1 Moments estimates
From the weak formulation (1.12) it is possible to derive a system of evolution equations for $M_{0},M_{1},M_{2}$ by choosing as test function $\left\{1,v,|v|^{2}\right\}$ respectively. It is important to note that the functions $\{1,v,|v|^{2}\}$ are not test functions in the sense of Definition 1 and we should consider suitable cut-offs in addition to obtain admissible test functions. However, assuming enough integrability for the solution $f$, the results obtained by relaxing the cut-offs are the same as the one we would obtain formally choosing $\{1,v,|v|^{2}\}$ as test functions. We refer, for instance, to Remark 6 in [48] for a detailed discussion concerning this technical issue. Therefore, to keep the exposition as simple as possible we will take $\{1,v,|v|^{2}\}$ as test functions in (1.12). Hence, using also the collision rule (1.10) leads to
$\begin{cases}\vskip 5.69054pt\displaystyle{\frac{\mathop{}\!\mathrm{d}M_{0}(t)}{\mathop{}\!\mathrm{d}t}=0},&\\ \vskip 5.69054pt\displaystyle{\frac{\mathop{}\!\mathrm{d}M_{1}(t)}{\mathop{}\!\mathrm{d}t}=aM_{0}(t)-(1+r)\int_{\mathbb{R}^{d}}f(t,dv)\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(v\cdot\omega)\omega\mathop{}\!\mathrm{d}\omega},&\\ \displaystyle{\frac{\mathop{}\!\mathrm{d}M_{2}(t)}{\mathop{}\!\mathrm{d}t}=2a\cdot M_{1}(t)-2(1-r^{2})\int_{\mathbb{R}^{d}}f(t,dv)\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(v\cdot\omega)^{2}\mathop{}\!\mathrm{d}\omega}.\end{cases}$ (3.1)
The system of equations (3.1) can be integrated exactly. This is the purpose of the next propositions.
@@
Let $d=2$, and let $a\in\mathbb{R}^{d}$ be a fixed vector. Assume that the restitution coefficient $r$ satisfies $0<r\leq 1$. Let $f\in C([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ be the unique weak solution to (1.9) given by Theorem 1.1. Suppose $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Then, we have that $M_{0}(t),M_{1}(t),M_{2}(t)\in L^{\infty}([0,+\infty))$.
@@
We start from (3.1) and we consider a rotation $R=R_{v}$ such that $N=Re_{1}$ (recalling that $N=v/|v|$) and introduce $\tilde{\omega}$ defined as $\omega=R\tilde{\omega}$. For $d=2$ we have $\tilde{\omega}=(\tilde{\omega}_{1},\tilde{\omega}_{2})=(\cos\theta,\sin\theta)$ with $\theta\in[0,2\pi]$. Therefore we have
@@
Therefore for we obtain the following equations:
@@
$M_{0}(t)$ $=M_{0}(0),$ (3.10)
$M_{1}(t)$ $=e^{-C_{1}t}M_{1}(0)+\frac{aM_{0}(0)}{C_{1}}\left(1-e^{-C_{1}t}\right),$ (3.11)
@@
Let $d=3$, and let $a\in\mathbb{R}^{d}$ be a fixed vector. Assume that the restitution coefficient $r$ satisfies $0<r\leq 1$. Let $f\in C([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ be the unique weak solution to (1.9) given by Theorem 1.1. Suppose $\{|v|,|v^{2}|\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Then, we have that $M_{0}(t),M_{1}(t),M_{2}(t)\in L^{\infty}([0,+\infty))$.
@@
$-(1+r)\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(v\cdot\omega)\omega\mathop{}\!\mathrm{d}\omega$ $=-2\pi(1+r)\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)v,$ (3.13)
$-(1-r^{2})\int_{\mathbb{S}^{d-1}}b(|N\cdot\omega|)(n\cdot\omega)^{2}\mathop{}\!\mathrm{d}\omega$ $=-2\pi(1-r^{2})\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)|v|^{2},$ (3.14)
@@
$M_{0}(t)$ $=M_{0}(0),$ (3.15)
$\frac{\mathop{}\!\mathrm{d}M_{1}(t)}{\mathop{}\!\mathrm{d}t}$ $=aM_{0}(0)-2\pi(1+r)\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)M_{1}(t),$ (3.16)
@@
Setting $\tilde{C}_{1}=2\pi(1+r)\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right),\tilde{C}_{2}=2\pi(1-r^{2})\left(\int_{-1}^{1}b(|x|)x^{2}\mathop{}\!\mathrm{d}x\right)$, we notice that $\tilde{C}_{1}-\tilde{C}_{2}>0$. Integrating we obtain:
$M_{0}(t)$ $=M_{0}(0),$ (3.18)
$M_{1}(t)$ $=e^{-\tilde{C}_{1}t}M_{1}(0)+\frac{aM_{0}(0)}{\tilde{C}_{1}}\left(1-e^{-\tilde{C}_{1}t}\right),$ (3.19)
@@ 3.2 Existence of a steady state
To prove existence of a steady state we will use the results of Proposition 3.1 and Proposition 3.2 which allow us to use Schauder’s fixed point theorem to prove the existence of a stationary solution to (1.9) in the sense of Definition 2. In fact, without loss of generality we choose $M_{0}(0)=1$ in (3.10) and (3.18), and we define the set:
$\mathscr{U}=\left\{f\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)\ /\ \int_{\mathbb{R}^{d}}f(\mathop{}\!\mathrm{d}v)=1,\int_{\mathbb{R}^{d}}|v|f(\mathop{}\!\mathrm{d}v)\leq aC_{0},\int_{\mathbb{R}^{d}}|v|^{2}f(\mathop{}\!\mathrm{d}v)\leq|a|^{2}\hat{C}_{0}\right\}$ (3.21)
@@
Let $d=2,3$, let $a\in\mathbb{R}^{d}$ be a fixed vector. Assume that the restitution coefficient $r$ satisfies $0<r<1$. Let $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ be such that $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ and let $f\in\mathcal{C}([0,+\infty),\mathscr{M}_{+}\left(\mathbb{R}^{d}\right))$ be the unique weak solution to (1.9) obtained in Theorem 1.1. Then, there exists a stationary solution $f_{\infty}$ of (1.12) in the sense of Definition 2, or equivalently, a fixed point for the adjoint $S^{*}(t)$ of the semigroup $S(t)$ defined by the duality formula (2.35).
@@
Due to Proposition 3.1 and Proposition 3.2, the set $\mathscr{U}$ is convex and closed in the weak-* topology of $\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. In addition, $\mathcal{U}$ is metrizable and hence sequentially compact and $\mathscr{U}$ is also weak-* compact. To see this, we observe that $\mathscr{U}$ is contained in the unit ball of $\mathscr{M}(\mathbb{R}^{d})$ and that the space $\mathcal{C}_{0}(\mathbb{R}^{d})$ is separable. Since $\mathscr{U}$ is weak-*-closed by the Banach-Alaoglu’s theorem (e.g., [19]), it follows that $\mathscr{U}$ is compact. We have that for any $h\geq 0$ $S^{*}(h)\mathscr{U}\subseteq\mathscr{U}$, hence the operator $S^{*}(h)$ is weak-* compact. Therefore we can apply Schauder’s fixed point theorem to prove the existence of $f_{*}^{(h)}$ such that $S^{*}(h)f_{*}^{(h)}=f_{*}^{(h)}$, and we have by the semigroup property that $S^{*}(mh)f_{*}^{(h)}=f_{*}^{(h)}$ for every $m\in\mathbb{N}$. We consider now a sequence $\{h_{k}\}_{k}$ such that $h_{k}\rightarrow 0$ and its corresponding sequence of fixed points $\big(f_{*}^{(h_{k})}\big)_{k}$. This sequence is compact in $\mathscr{U}$, since $\mathscr{U}$ itself is compact, and taking a sub-sequence if needed we have $f_{*}^{(h_{k})}\rightarrow f_{*}$ for some $f_{*}\in\mathscr{U}$ as $k\rightarrow+\infty$. For any $t>0$, there exists a sequence of integers such that $m_{k}h_{k}\rightarrow t$ as $k\rightarrow+\infty$. On the one hand, this yields $S^{*}(m_{k}h_{k})f_{*}^{(h_{k})}=S^{*}(h_{k})f_{*}^{(h_{k})}=f_{*}^{(h_{k})}\rightarrow f_{*}$, while on the other hand:
@@
By the weak-* continuity of the semigroup $S^{*}(t)$, it follows that the right-hand side converges to $S^{*}(t)f_{*}$, which gives:
@@
for every $t\geq 0$. Therefore $f_{*}$ is a fixed point for $S^{*}(t)$ with $t\geq 0$ arbitrary, and the proof is concluded. ∎
3.3 Uniqueness of the steady state
In this subsection we prove the uniqueness of the stationary solution obtained in Theorem 3.3. We will prove uniqueness expressing the stationary solution $f_{\infty}$ by expanding the equation $f_{\infty}=S^{*}(t)f_{\infty}$ by means of a Neumann series which we then prove to be convergent. This is a suitable adaptation of a well-known strategy, see for instance [6]. The main theorem of this subsection is the following.
@@ 3.3 Uniqueness of the steady state
Let $d=2$ or $3$, and let $a\in\mathbb{R}^{d}$ be a fixed vector. Assume that the restitution coefficient $r$ satisfies $0<r<1$. Let $f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$ be such that $\{|v|,|v|^{2}\}f_{0}\in\mathscr{M}_{+}\left(\mathbb{R}^{d}\right)$. Then, the stationary solution obtained in Theorem 3.3 is unique for every $t>0$.
@@
The semigroup $S(t)$ introduced in the proof of Theorem 1.1 acts on $\mathcal{C}_{0}(\mathbb{R}^{d})$ as
@@
where $F(t)$ is the semigroup generated by the operator $E[\varphi]=-a\cdot\partial_{v}\varphi-C_{b}\varphi$, where $C_{b}=\int_{\mathbb{S}^{d-1}}b\big(|N\cdot\omega|\big)\mathop{}\!\mathrm{d}\omega$, with domain $\mathscr{D}(T)$. The explicit action of $F(t)$ on $\mathcal{C}_{0}(\mathbb{R}^{d})$ is given by
$F(t)f=e^{-C_{b}t}f(v-at).$ (3.25)
@@
To compute the adjoint of $V(S(t))$ we consider $f\in\mathscr{M}(\mathbb{R}^{d}),\varphi\in\mathcal{C}_{0}(\mathbb{R}^{d})$, and we have that
@@
where $\langle\,$$\rangle$ denotes the duality bracket. Hence from (3.27) we can write
@@
where the series representation of $R(1,F^{*}(t))$ is absolutely convergent for any $t>0$. Hence (3.30) reads
@@
$\leq e^{-C_{b}nt}\sup_{t\geq 0}\|S(t)\|\|\mathscr{L}_{+}\|\int_{0}^{t}\|F^{*}(t-s)\|\mathop{}\!\mathrm{d}s$
$\leq e^{-C_{b}nt}C_{b}\int_{0}^{t}e^{-C_{b}(t-s)}\mathop{}\!\mathrm{d}s=e^{-C_{b}nt}(1-e^{-C_{b}t}).$ (3.33)
where we used that fact that $\|S(t)\|\leq 1$ for every $t\geq 0$ since $S(t)$ is a Markov semigroup and that $\|\mathscr{L}_{+}\|\leq C_{b}$. This gives
$\Bigg\|\sum_{n=0}^{+\infty}(F^{*}(t))^{n}(V(S(t)))^{*}\Bigg\|$ $\leq(1-e^{-C_{b}t})\sum_{n=0}^{+\infty}e^{-nC_{b}t}=(1-e^{-C_{b}t})\frac{1}{1-e^{-C_{b}t}}$ (3.34)
for any $t>0$ and shows that the series is converging. Since the Neumann series (3.32) identifies a unique element we conclude that the stationary solution $f_{\infty}$ is unique. ∎